Tom Roberts, CEO of lawnmower manufacturer Grassy Knowles, who has licensed software from the makers of Roomba in order to create the first self-directed, programmable lawnmower; he has branded this new mower the Goat. Knowles anticipates that fixed costs, including the plant and robotic assembly, will be $18 million for the first year. Projected sales for the first year are 90,000 units. Unit variable costs, including all mechanical and technical parts plus the licensing fee (contracted on a per-unit basis), add up to $1100 per mower. Using these parameters, project the price Tom would need to charge to breakeven. Show all calculations .
Answer #1

Let R be the price Tom needs to charge to break-even.

Fixed cost = $18000000

Variable cost per mower, v =$1100

Quantity = 90,000 units

At break-even point, TR = TC

R*Q = Fixed cost +(v*Q)

R*90000 = 18000000+(1100*90000)

Solving, R = $1300

The price Tom needs to charge to break-even = $1300 per unit.